OpenStudy (anonymous):

help to solve please. 51x=2x^2+72

OpenStudy (anonymous):

\[2x^2-51x+72=0\] I believe you have to use the quadratic equation again. \[x=-b \pm \sqrt{b^2-4ac}/2a \] \[ax^2+bx+c\] so a=2 b=-51 c=72

OpenStudy (anonymous):

I got x=24 or 1.5

OpenStudy (anonymous):

do you mind walking through it with me?

OpenStudy (anonymous):

yep, x=3/2 x=24

OpenStudy (anonymous):

\[-(-51)+\sqrt{(-51^2-4(2)(72)}/2(2)\]

OpenStudy (anonymous):

\[-(-51)-\sqrt{(-51^2-4(2)(72)}/2(2)\]

OpenStudy (anonymous):

Those are the equation you get when you plug a,b, and c into the quadratic equation.

OpenStudy (anonymous):

You have two equations because the quadratic equation has the opposite of b plus or minus the square root b squared minus four ac all divided by 2a.

OpenStudy (anonymous):

The first equation should give you a value of 24 and the second equation should give you a value of 1.5

OpenStudy (anonymous):

You can always check your answers as well by plugging them into the original equation. For instance \[2(24^2)-51(24)+72=0\]